OpenStudy (anonymous):

A family goes on vacation to a water park. During the drive to the park the weather caused the family to average 36 mph. After their vacation the weather improved, so they averaged 50 mph on their return trip. If the distance each way was equal, and the entire time driving was 8.6 hours. Find the distance one was to the water park.

OpenStudy (dumbcow):

t1 + t2 = 8.6 36*t1 = 50*t2 --> t1 = (25/18)*t2 (25/18)*t2 + t2 = 8.6 t2(25/18 + 1) = 8.6 t2 = 8.6/2.388 = 3.6 d = 50*3.6 = 180 miles

OpenStudy (anonymous):

Wonderful, thank you!

OpenStudy (anonymous):

Yeah its 180 miles.